Suppose a firms demand curve is P = 100 - Q. Find MR when Q = 10. Please show work
Answer #1

The demand function is given as:

P = 100 - Q

The total revenue is equal to the product of price and quantity. So,

Total Revenue = PQ = 100Q - Q²

The marginal revenue is calculated by differentiating the total revenue function with respect to quantity:

Marginal Revenue = 100 - 2Q

The marginal revenue when Q = 10 is 100 - (2)(10) = 100 - 20 = 80.

So, the marginal revenue equals 80 when the quantity is 10.